# Exploring Families’ Use and Mediation of Generative AI: A Multi-User Perspective

### v2

@@ Abstract.
Applications of Generative AI (GenAI), such as ChatGPT, have gained popularity among the public due to their ease of access, use, and support of educational and creative activities. Despite these benefits, GenAI poses unique risks for families, such as lacking sufficient safeguards tailored to protect children under 16 years of age and not offering parental control features. This study explores families’ use and co-use of GenAI, the perceived risks and opportunities of ChatGPT, and how parents mediate their children’s use of GenAI. Through semi-structured interviews with 12 families, we identified ways families used and mediated GenAI and factors that influenced parents’ GenAI mediation strategies. We contextualize our findings with a modified model of family mediation strategies, drawing from previous family media and mediation frameworks. We provide insights for future research on family-GenAI interactions and highlight the need for more robust protective measures on GenAI platforms for families.
Figure 1. Summary of Observed Family-GenAI Profiles: We interviewed 12 families about their use and co-use of text-based generative AI platforms (i.e., ChatGPT). We contextualized our findings into six observed Family-GenAI mediation profiles. The six profiles (i.e., skeptical, cautious, hands-on, together, trusting, and independent families) demonstrate different types of Family-GenAI use, with respect to the following three dimensions: parents’ control over children’s ChatGPT usage, families’ trust in ChatGPT’s safeguards and information quality, and families’ frequency of co-use.
[…]
A table depicting the six types of family-AI use patterns identified in the present paper. The ”Skeptical” Family is labeled as ‘sets protective measures for use’ and ‘High control, low trust, medium co-use’ is depicted as a caregiver saying, “I’m putting in some guardrails”, as their child observes beside them. The ”Cautious” Family is labeled as ‘Sets exploratory rules for monitoring’ and ‘high control, high trust, low co-use’ is depicted as a child using a laptop while their parent faces away, but peers at the laptop screen behind their shoulder and says, “ I keep an eye on them.” The ”Hands-On” Family is labeled as ‘actively screens child’s use’ and ‘High control, medium trust, medium co-use’ is depicted as a child using a laptop as their parent stands behind them to observe the screen and says, “I use my personal judgment.” The ”Use Together” Family is labeled as ‘Actively co-uses together’ and ‘High control, medium trust, high co-use’ and is depicted as a child and parent using a laptop together as the parent says, “We try to share accounts.” The ”Trusting” Family is labeled as ‘Transfers responsibility to the platform’ and ‘Medium control, high trust, medium co-use’ and is depicted as a child and parent back-to-back while using a cellphone and a laptop, respectively, but looking at each other as the parent says, “There are limits built in.” The ”Independent” Family is labeled as ‘Relaxed on rules’ and ‘Low control, medium trust, low co-use’ and is depicted as a child and parent back-to-back and looking away from one another while using a laptop and a cellphone, respectively, with the parent saying, “I can’t always be there for them.”
@@ 1. Introduction
Imagine a family that is looking for a third player for their favorite card game and teaching the rules to ChatGPT to play with them. Later, during a family debate, the parent suggests, “Let’s ask ChatGPT to settle this!” Scenarios like these are becoming increasingly common as generative AI (GenAI) and large language model (LLM)-based chatbots (e.g., ChatGPT, Gemini, Claude) become integral to daily life (Bommasani et al., 2021). However, these interactions also raise concerns for parents, ranging from fears such as, “What if my child uses ChatGPT to cheat on homework?” to more speculative worries like, “I should be kind to AI in case it takes over the world one day.” Understanding how GenAI is beginning to influence family routines and activities requires examining how families harness the potential of this emerging technology while navigating its associated risks.
@@
Today’s parents may face struggles with deciding whether and how to integrate GenAI into their lives, much like they did with previous technologies such as the Internet, search engines, and smartphones (Park and Lim, 2020; Long et al., 2022). GenAI and LLM-based chatbots offer significant potential to support children’s creative and educational pursuits (Jeon and Lee, 2023; Krause et al., 2024; Druga et al., 2019) because they can generate content and information in conversational, adaptable language (Akhtar, 2024). While families may apply lessons from online safety with older technologies (e.g., Internet, TV, social media) to GenAI, their level of familiarity with GenAI may limit their ability to address its unique risks (Yu et al., 2024). Parents often worry about inappropriate content and interactions online, and emerging research highlights new risks, such as emotional dependence on GenAI, misinformation from hallucinations, and unchecked trust in GenAI (Yu et al., 2024; Solyst et al., 2024). Our study aims to explore how families navigate these risks and opportunities.
@@
Parents use various mediation strategies to reduce the risks associated with their children’s use of media and technology (Valkenburg et al., 1999). These strategies depend on the technology’s features (Nikken and Jansz, 2014), parents’ perceptions of it (Clark, 2011), and family factors like the child’s age and parental attitudes (Rasmussen et al., 2016; Chen and Shi, 2019). For instance, parents may overlook data privacy risks if they are unaware that GenAI platforms retain inputted prompts and responses unless manually deleted. Few studies have examined parents’ GenAI mediation strategies. While some studies show parents attempting to manage their children’s use of GenAI, they also highlight gaps in parental knowledge that limit their ability to address the technology’s risks (Yu et al., 2024; Abel and Magnusson, 2024). Although no technology is risk-free, effective use of GenAI can enhance family interactions (Kildare and Middlemiss, 2017), improve children’s social connections (Naslund et al., 2020), and support academic learning (Ochoa and Reich, 2020). Overall, this study aims to explore how and why parents mediate their children’s use of GenAI, focusing on the risks, opportunities, and the role of parental perceptions and family contexts. To our knowledge, little existing research addresses these perspectives in-depth.
We pose the following research questions:
  • RQ1: What mediation strategies do parents use for GenAI?
  • RQ2: What benefits or risks do parents perceive in their child’s use of GenAI?
  • RQ3: What contextual factors shape families’ perceptions of GenAI and parents’ GenAI mediation strategies?
@@
To address these questions, we take a multi-user perspective, focusing on shared family use of GenAI. We conducted exploratory semi-structured interviews with 12 families that consisted of at least one child with experience sharing a ChatGPT account with a parent. We analyze our findings using existing parental media mediation frameworks and propose a modified model for family use of GenAI. Our work aims to inform efforts to evaluate and improve family safety on GenAI platforms.
2. Related Work and Background
@@ 2.1. Children, Families, and GenAI
Since its release in 2022, ChatGPT and other GenAI tools have become increasingly prevalent in children’s lives. Surveys indicate that there are many users under 18 (Sidoti and Gottfried, 2023; Research, 2024; Gottfried and Jeffrey, 2023; Media, 2023, 2024), with one study reporting children as young as five interacting with ChatGPT (Quan et al., 2024). It is important to acknowledge that GenAI use exposes children to risks such as misinformation, data misuse (Wach et al., 2023), and more alarming concerns like deepfake pornography and harmful interactions (NBC News, Angela Yang, 2025; New York Post, 2025). As a result, an increasing amount of research has focused on Child-Centered AI, and GenAI privacy and safety  (Atabey et al., 2024; Wang et al., 2023; Honauer and Frauenberger, 2024; Zhang et al., 2022; Wang et al., 2022; Druga et al., 2022; Ragone et al., 2024; Kurian, 2024; Yu et al., 2024; Solyst et al., 2024). At home, parents may try to manage potential risks of GenAI use by employing various mediation strategies (Yu et al., 2024; Abel and Magnusson, 2024). However, if parents are unfamiliar with GenAI, they may find it more difficult to manage their children’s GenAI use  (Yu et al., 2024), and are likely to encounter limited ways to control use in-platform. Unlike streaming or social media platforms with tailored protections for minors (Netflix, [n. d.]; YouTube, [n. d.]), many GenAI platforms currently lack child-specific safeguards. This is partly because many of the platforms require users to be 13 or older (OpenAI, [n. d.]), though research and anecdotal evidence suggest limited enforcement of this rule (Quan et al., 2024; Research, 2024). At the policy level, some researchers argue that overarching policies meant to protect children online, such as the Kids Online Safety Act and COPPA (U.S. Congress, [n. d.]; Congress, 1998), are ineffective or potentially harmful to young people (Angel and Boyd, 2024; Lwin et al., 2008). There is an urgent need for both platform-level protections and family-centered tools to better safeguard children in the evolving GenAI landscape.
@@ 2.1.1. Technology at Home
Along with addressing parental concerns, it is also important to acknowledge parents’ motivations for providing access to GenAI platforms at home. There is limited mediation literature about reasons for GenAI use in households with children, but work in adjacent fields provides valuable insights into what parents value or consider when introducing new technology. While most forms of technology present some sort of risk to children, both children and parents also report benefits from their use. Various devices and media can support family life and child development. For example, distanced family members can be connected through social media and devices (Neustaedter et al., 2015). High-quality interactive and non-interactive media, including device applications and television, have been linked to long-term educational outcomes and short-term promotion of play in young children (Marsh et al., 2018; Kearney and Levine, 2019). Families have also reported using media together (e.g., video games and Netflix) as a tool to bond with one another (Sobel et al., 2017; Wang et al., 2018). Preliminary research on GenAI suggests that some parents believe GenAI can positively impact their children’s education and have used it to facilitate their children’s learning (Quan et al., 2024).
@@ 2.1.2. Evaluating Families’ Technology Use
Families’ knowledge, needs, and goals shape their home technologies and related rules. According to a national survey by  Rideout and Robb (2020) on children’s screen media use in the US, parents from historically marginalized racial and socioeconomic backgrounds are more likely to view screen media as educational. Additionally, parents’ mediation strategies have been correlated with parents’ attitudes toward specific media (Clark, 2011). These attitudes and perceptions may be motivated by the amount of knowledge a parent has obtained about the media based on their personal experiences and the external educational resources they may access to (Nikken and Jansz, 2014), as well as their personal and their children’s behaviors  (Milford et al., 2022). At this time, there is little available work to draw parallel conclusions about GenAI, but at least one study has noted that parents’ occupations may contribute to whether they use GenAI tools for work or entertainment (Abel and Magnusson, 2024).
@@ 2.2. Adapting Parental Mediation Strategies For Emerging Technologies
In their seminal work, Valkenburg et al. (1999) identified three types of parental mediation for traditional media (i.e., television): restrictive mediation (limiting access or use), instructive or active mediation (parent-child conversations about media content), and co-use (parents and children using media together). These strategies have remained relevant across technologies (Nikken and Jansz, 2006, 2014; Yu et al., 2024), but their definitions have evolved to address new aspects of emerging media. Scholars have also identified additional strategies tailored to modern technologies, such as monitoring internet use (Nikken and Jansz, 2014), preparing hybrid digital toys for children (Yu et al., 2021), and learning from children during co-use (Clark, 2011). The specific features of technologies often determine the suitability of certain strategies; for example, parents tend to favor co-use and instructive mediation over restrictive approaches for digital educational programs (Yu et al., 2021). Over time, parental strategies for the same technology may also shift as perceptions and societal norms evolve. While  Mascheroni and Ólafsson (2014) suggested that monitoring children’s mobile phone use was difficult because devices were meant to be private,  Ren and Zhu (2022) later highlighted monitoring as a common mediation strategy for managing mobile phone use.
@@
Although limited, emerging research on GenAI mediation strategies suggests, parents rely on both established and new approaches. Both  Yu et al. (2024) and  Abel and Magnusson (2024) found evidence of parents using instructive, restrictive, and co-use mediation strategies to manage GenAI use. However, parents’ limited understanding of GenAI’s mechanisms often prevented them from addressing risks specific to the technology (e.g., not understanding the implications of generative chatbots). Beyond traditional strategies, some parents have adopted approaches to prepare for their child’s use of GenAI. For instance,  Abel and Magnusson (2024) observed parents educating themselves and their children about GenAI in anticipation of future use, while  Quan et al. (2024) found parents occasionally encouraged their children to engage with GenAI. Given the gaps in understanding how family dynamics influence the adoption and mediation of GenAI, this study aims to explore parents’ perceptions of GenAI’s risks and benefits and to identify family characteristics that shape how GenAI is (co-)used and mediated at home.
@@
In this work, we aimed to explore how families with children navigate the access and use of a shared ChatGPT account via semi-structured interviews. Our work differs from prior work on children’s or individuals’ use of GenAI by focusing on families with children to reveal unique insights about their various experiences within this domain. We draw insights from relevant family media mediation literature and propose an adapted version of prior theoretical models (e.g., Nikken and Jansz (2006)) to describe family-GenAI mediation and interactions. Based on our findings, we also identify several risks and opportunities unique to GenAI that families may encounter when using GenAI platforms. Throughout this investigation, we take a “multi-user perspective” to capture both parents’ and children’s experiences using GenAI.
@@ 3.1. Participants and Recruitment
We recruited 12 U.S.-based families via paid advertisement on social media platforms (Meta Ads [^1] https://business.meta.com/?locale=en_US) and e-mail lists. We distributed our pre-screening survey in our advertisements and contacted eligible respondents who (1) was a parent of at least one child between the ages of 8 and 18 who had shared a ChatGPT account with them and was willing to take part in the study, (2) spoke English, and (3) resided in the United States. All interviews were conducted over Zoom. Participants over 18 (i.e., the parents in each family) signed a consent form before participating in the study. Minors were verbally assented at the start of each Zoom call. Parents were also asked to complete a demographics form at the end of the call. Families were compensated \$25 after the interview. Of the 12 families, two had two children present during the interview, and two had two parents present. All other family interviews consisted of one child and one parent. Fourteen parents total participated in the study (10 mothers and four fathers). 42% of the participating children were girls, and 58% were boys. 71% of the participating parents held at least a Bachelor’s degree, and all participating parents reported that they were at least “slightly familiar with new technologies.” Detailed demographics of the participating families can be found in Table 1.
Table 1. Demographics
@@
  • Columns 1 & 8 are family demographics, columns 2-5 are child demographics, and columns 6-7 are parent demographics. All demographics were self-reported by parents. All occupation information addressed in the later passage was self-disclosed during the interview; we did not explicitly collect any information about location or occupation information to maintain families’ privacy.

### v3

@@ Abstract.
Generative AI (GenAI) platforms, such as ChatGPT, have gained popularity among the public due to their ease of access, use, and support of educational and creative activities. Despite these benefits, GenAI poses unique risks for families, such as lacking sufficient safeguards tailored to protect children under 13 years of age and not offering parental control features. This study explores how parents mediate their children’s use of GenAI and the factors/processes that shape this mediation. Through analyzing semi-structured interviews with 12 families, we identified ways in which families used and mediated GenAI and factors that influenced parents’ GenAI mediation strategies. We contextualize our findings with a modified model of family mediation strategies, drawing from previous family media and mediation frameworks. We provide insights for future research on Family-GenAI interactions and highlight the need for more robust protective measures on GenAI platforms for families.
Figure 1. Summary of Observed Family-GenAI Profiles: We interviewed 12 families about their use and co-use of text-based generative AI platforms (i.e., ChatGPT). We contextualized our findings into six observed Family-GenAI mediation profiles. The six profiles (i.e., skeptical, cautious, hands-on, together, trusting, and independent families) demonstrate different types of Family-GenAI use, with respect to the following three dimensions: parents’ control over children’s ChatGPT usage, families’ trust in ChatGPT’s safeguards and information quality, and families’ frequency of co-use.:
@@ 1. Introduction
Imagine a family that is looking for a third player for their favorite card game and teaching the rules to ChatGPT to play with them. Later, during a family debate, the parent suggests, “Let’s ask ChatGPT to settle this!” Scenarios like these are becoming increasingly common as generative AI (GenAI) and large language model (LLM)-based chatbots (e.g., ChatGPT, Gemini, Claude) become integral to daily life (bommasani2021opportunities). However, these interactions also raise concerns for parents, ranging from fears such as, “What if my child uses ChatGPT to cheat on homework?” to more speculative worries like, “I should be kind to AI in case it takes over the world one day.” Understanding how GenAI is beginning to influence family routines and activities requires examining how families harness the potential of this emerging technology while navigating its associated risks.
@@
Today’s parents may face struggles with deciding whether to integrate and how to manage GenAI into their families, much like they did with previous technologies such as the Internet, search engines, and smartphones (park2020investigating; long2022family). Although no technology is risk-free, effective use of GenAI can enhance family interactions (KildareMiddlemiss2017), improve children’s social connections (naslund2020social), and support academic learning (ochoa2020parents). For example, GenAI and LLM-based chatbots offer significant opportunities to support children’s creative and educational pursuits (jeon2023large; krause2024evolution; druga2019inclusive) because they can generate content and information in conversational, adaptable language (akhtar2024unveiling). Still, parents often worry about inappropriate content and interactions online, and emerging research highlights new risks, such as emotional dependence on GenAI, misinformation from hallucinations, and unchecked trust in GenAI (yu2024exploring; solyst2024children). Our study explores how families navigate these opportunities and risks.
@@
Before GenAI, parents used various mediation strategies to reduce the risks associated with their children’s use of media and technology (valkenburg1999developing). These strategies depend on the technology’s features (nikken2014developing), parents’ perceptions of the technology (clark2011parental), and familial characteristics like the child’s age and parental attitudes (rasmussen2016predicting; chen2019reducing). While families may apply lessons from online safety with older technologies (e.g., Internet, TV, and social media) to GenAI, their level of familiarity with GenAI may limit their ability to address its unique risks (yu2024exploring). For instance, parents may overlook data privacy risks if they are unaware that GenAI platforms retain inputted prompts and responses unless manually deleted. Few studies have examined parents’ GenAI mediation strategies. Some studies show parents attempting to manage their children’s use of GenAI. They also highlight gaps in parental knowledge that limit their ability to address the technology’s risks (yu2024exploring; abel2024playdates). However, there is little research on how and why parents mediate their children’s use of GenAI, focusing on risks, opportunities, and the role of parental perceptions and family contexts.
@@
To bridge this gap, we take a multi-user perspective, focusing on shared family use of GenAI. We analyzed semi-structured interviews with 12 families that consisted of at least one child with experience sharing a ChatGPT account with a parent. We analyze our findings using existing parental media mediation frameworks and propose a modified model for family use of GenAI. Our work aims to inform efforts to evaluate and improve family safety on GenAI platforms. Throughout this investigation, we take a “multi-user perspective” to capture both parents’ and children’s experiences using GenAI.
Our contributions include: (1) We present emerging themes from parent and child co-use of ChatGPT. (2) Informed by the emerging themes, we propose a model to conceptualize the factors affecting the decision-making in family GenAI mediation. (3) We describe family interactions around GenAI through the lens of the developed model. (4) We propose design implications for family-centric GenAI mediation from the views of multiple stakeholders.
2. Background & Related Work
@@ 2.1. Children, Families, and GenAI
Since its release in 2022, ChatGPT and other GenAI tools have become increasingly prevalent in children’s lives. Surveys indicate that there are many users under 18 (sidoti20231; 4h2024; Gottfried_2023; commonsense2023ai; commonsense2024ai), with one study reporting children as young as five interacting with ChatGPT (quan2024young). It is important to acknowledge that GenAI use exposes children to risks such as misinformation, data misuse (wach2023dark), and more alarming concerns like deepfake pornography and harmful interactions (nbcnews_characterai_2025; nypost_deepfake_2025). As a result, an increasing amount of research has focused on Child-Centered AI, and GenAI privacy and safety  (atabey2024second; wang2023child; honauer2024exploring; zhang2022storybuddy; wang2022informing; druga2022family; ragone2024child; kurian2024no; yu2024exploring; solyst2024children). At home, parents may try to manage potential risks of GenAI use by employing various mediation strategies (yu2024exploring; abel2024playdates). However, if parents are unfamiliar with GenAI, they may find it more difficult to manage their children’s GenAI use  (yu2024exploring), and are likely to encounter limited ways to control use in-platform. Unlike streaming or social media platforms with tailored protections for minors (Netflixcontrols2022; YouTubesafety2024), many GenAI platforms currently lack child-specific safeguards. This is partly because many of the platforms require users to be 13 or older (OpenAIterms2024), though research and anecdotal evidence suggest limited enforcement of this rule (quan2024young; 4h2024). At the policy level, some researchers argue that overarching policies meant to protect children online, such as the Kids Online Safety Act and COPPA (KidsOnlineSafetyAct; coppa1998), are ineffective or potentially harmful to young people (angel2024techno; lwin2008protecting). There is an urgent need for both platform-level protections and family-centered tools to better safeguard children in the evolving GenAI landscape.
@@ 2.1.1. Technology at Home
Along with addressing parental concerns, it is also important to acknowledge parents’ motivations for providing access to GenAI platforms at home. While there is limited literature regarding the motivation for GenAI use in households with children, existing work does highlight positive developmental and learning outcomes (sun2024; quan2024young; kumar2024; neshaei2025; zhang2025qualitative). Additionally, works in adjacent fields provide valuable insights into what parents value or consider when introducing new technology. While most forms of technology present some sort of risk to children, both children and parents also report benefits from their use. For example, families may use technologies such as digital games, VR, and social media to bridge intergenerational (pais2024; du2025; Rubin2025) and geographical divides  (neustaedter2015sharing; yuan2024). High-quality interactive and non-interactive media, including device applications and television, have been linked to long-term educational outcomes and short-term promotion of play in young children (Marsh_Plowman_Yamada-Rice_Bishop_Lahmar_Scott_2018; Kearney_Levine_2019). Families have reported using media together (e.g., video games and Netflix) as a tool to bond with one another (sobel2017wasn; wang2018families). Such bonding is also identified as a primary motivator in Zhang et al.’s work studying families’ use and gratifications of ChatGPT (zhang2025qualitative).
@@ 2.1.2. Evaluating Families’ Technology Use
Families’ knowledge, needs, and goals shape their home technologies and related rules. According to a national survey by  Rideout_Robb_2020 on children’s screen media use in the US, parents from historically marginalized racial and socioeconomic backgrounds are more likely to view screen media as educational. Additionally, parents’ mediation strategies have been correlated with parents’ attitudes toward specific media (clark2011parental). These attitudes and perceptions may be motivated by the amount of knowledge a parent has obtained about the media based on their personal experiences and the external educational resources they may access to (nikken2014developing), as well as their personal and their children’s behaviors  (milford2022initial). At this time, there is little available work to draw parallel conclusions about GenAI, but at least one study has noted that parents’ occupations may contribute to whether they use GenAI tools for work or entertainment (abel2024playdates).
@@ 2.2. Adapting Parental Mediation Strategies For Emerging Technologies
In their seminal work, valkenburg1999developing identified three types of parental mediation for traditional media (i.e., television): restrictive mediation (limiting access or use), instructive or active mediation (parent-child conversations about media content), and co-use (parents and children using media together). These strategies have remained relevant across technologies (nikken2006parental; nikken2014developing; yu2024exploring), but their definitions have evolved to address new aspects of emerging media. Scholars have also identified additional strategies tailored to modern technologies, such as monitoring internet use (nikken2014developing), preparing hybrid digital toys for children (yu2021parental), and learning from children during co-use (clark2011parental). The specific features of technologies often determine the suitability of certain strategies; for example, parents tend to favor co-use and instructive mediation over restrictive approaches for digital educational programs (yu2021parental). Over time, parental strategies for the same technology may also shift as perceptions and societal norms evolve. While  mascheroni2014risks suggested that monitoring children’s mobile phone use was difficult because devices were meant to be private,  ren2022parental later highlighted monitoring as a common mediation strategy for managing mobile phone use.
@@
Although limited, emerging research on GenAI mediation strategies suggests, parents rely on both established and new approaches. Both  yu2024exploring and  abel2024playdates found evidence of parents using instructive, restrictive, and co-use mediation strategies to manage GenAI use. However, parents’ limited understanding of GenAI’s mechanisms often prevented them from addressing risks specific to the technology (e.g., not understanding the implications of generative chatbots). Beyond traditional strategies, some parents have adopted approaches to prepare for their child’s eventual use of GenAI. For instance,  abel2024playdates observed parents educating themselves and their children about GenAI in anticipation of future use, while  quan2024young found parents occasionally encouraged their children to engage with GenAI. Given the gaps in understanding how family dynamics influence the adoption and mediation of GenAI, this study aims to explore parents’ perceptions of GenAI’s risks and benefits and to identify family characteristics that shape how GenAI is (co-)used and mediated at home.
@@
In this work, we study how families with children navigate the use of a shared ChatGPT account, and we delve deeper into their mediation strategies. Our work differs from prior work on children’s or individuals’ use of GenAI by focusing on families with children to reveal unique insights about their various experiences within this domain. We draw insights from relevant family media mediation literature and propose an adapted version of prior theoretical models (e.g., nikken2006parental) to describe family-GenAI mediation and interactions.
@@ 3.1. Participants and Recruitment
In this work, we leverage the dataset collected by Zhang et al. (zhang2025qualitative). They recruited 12 U.S.-based families via paid advertisement on social media platforms (Meta Ads [^1] https://business.meta.com/?locale=en_US) and e-mail lists. The inclusion criteria for that data included respondents who (1) were a parent of at least one child between the ages of 8 and 18, (2) shared a ChatGPT account with family members, (3) spoke English, and (4) resided in the United States. All interviews were conducted over Zoom and followed proper ethical procedures in terms of obtaining IRB approval, consent from adult parents, and assent from children. Families were compensated \$25 after the interview.
Of the 12 families, two had two children present during the interview, and two had two parents present. All other family interviews consisted of one child and one parent. Fourteen parents total participated in the study (10 mothers and four fathers). 42% of the participating children were girls, and 58% were boys. 71% of the participating parents held at least a Bachelor’s degree, and all participating parents reported that they were at least “slightly familiar with new technologies.” Detailed demographics of the participating families can be found in Table 1.
Table 1. Demographics of the participants in the study conducted by Zhang et al. (zhang2025qualitative)
@@
  • Columns 1 & 8 are family demographics, columns 2-5 are child demographics, and columns 6-7 are parent demographics. All demographics were self-reported by parents. All occupation information addressed in the later passage was self-disclosed during the interview.